OpenStudy (anonymous):

(a) Solve 3cosθ+2=0 for 0°≤θ≤360°. --->132°, 228° (b) Hence solve 3cos^(2)θ+2cosθ=0 for 0°≤θ≤180°. @hartnn

OpenStudy (anonymous):

i don't know how to do part b.

hartnn (hartnn):

3c^2+2c = 0 c (3c+2) = 0 c=0 or 3c +2=0

hartnn (hartnn):

so what angle makes cos theta = 0 ?

OpenStudy (anonymous):

90, 270(rej)

OpenStudy (anonymous):

∴90

hartnn (hartnn):

one will be rejected from 132°, 228° too

OpenStudy (anonymous):

228

hartnn (hartnn):

so u'll have 2 angles

OpenStudy (anonymous):

oh i see. haha i am so stupid...